# Scalable unsupervised feature selection via weight stability

### v1

@@ Abstract
Unsupervised feature selection is critical for improving clustering performance in high-dimensional data, where irrelevant features can obscure meaningful structure. In this work, we introduce the Minkowski weighted $k$-means++, a novel initialisation strategy for the Minkowski Weighted $k$-means. Our initialisation selects centroids probabilistically using feature relevance estimates derived from the data itself. Building on this, we propose two new feature selection algorithms, FS-MWK++, which aggregates feature weights across a range of Minkowski exponents to identify stable and informative features, and SFS-MWK++, a scalable variant based on subsampling. We support our approach with a theoretical guarantee under mild assumptions and extensive experiments showing that our methods consistently outperform existing alternatives.
@@ 3.2 Feature Selection With MWK++
We now formally justify our use of weight stability across different values of $p$ as the basis for FS-MWK++ by presenting a theoretical guarantee. We begin with a short definition and a lemma.
@@
Let $v$ be a noise feature, and suppose that all such features are drawn independently from the same distribution and are uncorrelated with cluster structure. If there exists at least one relevant feature, then for any $p>1$,
$w_{lv}^{(p)}<\frac{1}{m}.$
@@
Suppose $u^{*}$ is a relevant feature. Since relevant features are aligned with cluster structure, we have that
$D_{lu^{*}}^{(p)}<D_{lv}^{(p)}.$
Let $r=\left(\frac{D_{lv}^{(p)}}{D_{lu^{*}}^{(p)}}\right)^{\frac{1}{p-1}}>1$. Now consider the sum in the denominator of Equation (3),
$\sum_{u=1}^{m}\left(\frac{D_{lv}^{(p)}}{D_{lu}^{(p)}}\right)^{\frac{1}{p-1}}=r+\sum_{u\neq u^{*}}\left(\frac{D_{lv}^{(p)}}{D_{lu}^{(p)}}\right)^{\frac{1}{p-1}}.$
Since all noise features have dispersions similar to $D_{lv}^{(p)}$, for each irrelevant $u$,
$\left(\frac{D_{lv}^{(p)}}{D_{lu}^{(p)}}\right)^{\frac{1}{p-1}}\approx 1\Rightarrow\sum_{u=1}^{m}\left(\frac{D_{lv}^{(p)}}{D_{lu}^{(p)}}\right)^{\frac{1}{p-1}}\approx(m-1)+r>m\Rightarrow w_{lv}^{(p)}<\frac{1}{m}.$
This lemma allows us to show that, under mild assumptions, relevant features have a higher median weight than noise features when aggregated across clusters and distance exponents. The next theorem formalises this insight.
@@
Let $P\neq\emptyset$ be a finite set of valid Minkowski exponents. If $u^{*}$ is a relevant feature for at least half of the clusters and $v$ is a noise feature drawn independently from a common distribution and uncorrelated with cluster structure, then
$\mathrm{median}\left(\{w_{lu^{*}}^{(p)}:p\in P,\,l=1,\dots,k\}\right)>\mathrm{median}\left(\{w_{lv}^{(p)}:p\in P,\,l=1,\dots,k\}\right).$
@@
From Lemma 1, we have that for any $p>1$
$w_{lv}^{(p)}\in\{w_{lv}^{(p)}:p\in P,\,l=1,\dots,k\}\Rightarrow w_{lv}^{(p)}<\frac{1}{m}.$
Hence, $\mathrm{median}\left(\{w_{lv}^{(p)}:p\in P,\,l=1,\dots,k\}\right)<\frac{1}{m}$.
Now consider the set $\{w_{lu^{*}}^{(p)}:p\in P,\,l=1,\dots,k\}$. Observe that Equation (3) is strictly convex for $p>1$, so
$w_{lu^{*}}^{(p)}=\frac{1}{m}\iff\sum_{u=1}^{m}\left[\frac{D_{lu^{*}}}{D_{lu}}\right]^{\frac{1}{p-1}}=m\iff\frac{D_{lu^{*}}}{D_{lu}}=1\text{ for all }u=1,\ldots,m.$
However, the existence of $v$ shows $X$ has at least one noise feature. Thus, the above cannot happen and by consequence $w_{lu^{*}}^{(p)}\neq\frac{1}{m}$. Recall that (3) ensures the weights for a particular cluster $S_{l}$ must sum to one, and that $w_{lv}^{(p)}<\frac{1}{m}$. Hence, the difference $\frac{1}{m}-w_{lv}^{(p)}>0$ is spread over any relevant feature, and $w_{lu^{*}}^{(p)}>\frac{1}{m}$. We have, by hypothesis, that $u^{*}$ is relevant for at least half of the clusters. Therefore, $\mathrm{median}\left(\{w_{lu^{*}}^{(p)}:p\in P,\,l=1,\dots,k\}\right)>\frac{1}{m}$. ∎
@@ 4 Experiments setting
We divide our experiments into two main sets to separately evaluate (i) the clustering performance of the proposed MWK++ initialisation method, and (ii) the effectiveness of FS-MWK++ and SFS-MWK++ for unsupervised feature selection.
@@ 6 Conclusion
Overall, our results highlight the potential of clustering-driven feature weighting as a robust foundation for unsupervised learning tasks. Future work will explore theoretical guarantees for weight stability under different $p$ values, and extend our methods to more general clustering frameworks beyond MWK.

### v2

@@ Abstract
Unsupervised feature selection is critical for improving clustering performance in high-dimensional data, where irrelevant features can obscure meaningful structure. In this work, we introduce the Minkowski weighted $k$-means++, a novel initialisation strategy for the Minkowski Weighted $k$-means. Our initialisation selects centroids probabilistically using feature relevance estimates derived from the data itself. Building on this, we propose two new feature selection algorithms, FS-MWK++, which aggregates feature weights across a range of Minkowski exponents to identify stable and informative features, and SFS-MWK++, a scalable variant based on subsampling. We support our approach with a theoretical guarantee under mild assumptions and extensive experiments showing that our methods consistently outperform existing alternatives. Our software can be found at https://github.com/xzhang4-ops1/FSMWK.
@@ 3.2 Feature Selection With MWK++
To establish theoretical guarantees for feature selection, we identify a sufficient condition under which a feature consistently receives high weight across multiple clustering configurations. This condition provides the foundation for analysing when FS-MWK++ is guaranteed to select a relevant feature.
@@
Let $t$ be a noise feature, and suppose that all such features are drawn independently from the same distribution and are uncorrelated with cluster structure. If there exists at least one relevant feature, then for any $p>1$,
$w_{lt}^{(p)}<\frac{1}{m}.$
@@
Let $\mathcal{W}$ be the set of noise features. Then, for a particular cluster $S_{l}$ the denominator of Equation (3) becomes
$\sum_{u=1}^{m}\left(\frac{D_{lt}^{(p)}}{D_{lu}^{(p)}}\right)^{\frac{1}{p-1}}=\sum_{u\in\mathcal{W}}\left(\frac{D_{lt}^{(p)}}{D_{lu}^{(p)}}\right)^{\frac{1}{p-1}}+\sum_{u\not\in\mathcal{W}}\left(\frac{D_{lt}^{(p)}}{D_{lu}^{(p)}}\right)^{\frac{1}{p-1}}.$
Given all noise features have similar dispersions,
$\frac{D_{lt}^{(p)}}{D_{lu}^{(p)}}\approx 1\text{ for all }u\in\mathcal{W}\Rightarrow\sum_{u\in\mathcal{W}}\left(\frac{D_{lt}^{(p)}}{D_{lu}^{(p)}}\right)^{\frac{1}{p-1}}\approx|\mathcal{W}|.$
Noise features are uncorrelated with cluster structure. Hence, their dispersion will be higher than that of a relevant feature. Thus,
$\frac{D_{lt}^{(p)}}{D_{lu}^{(p)}}>1\text{ for all }u\not\in\mathcal{W}\Rightarrow\sum_{u\not\in\mathcal{W}}\left(\frac{D_{lt}^{(p)}}{D_{lu}^{(p)}}\right)^{\frac{1}{p-1}}>m-|\mathcal{W}|.$
Finally,
$\sum_{u=1}^{m}\left(\frac{D_{lt}^{(p)}}{D_{lu}^{(p)}}\right)^{\frac{1}{p-1}}>m\Rightarrow w_{lt}^{(p)}<\frac{1}{m}.$
Lemma 2.
Let $X$ contain both relevant and noise features. If each noise feature is drawn independently from a common distribution, then at least one feature will have a weight strictly higher than $\frac{1}{m}$.
Proof.
From Lemma 1 we have that if $t$ is a noise feature, then $w_{lt}^{(p)}<\frac{1}{m}$ for any $p>1$ and cluster $S_{l}$. Since the feature weights for each cluster $S_{l}$ sum to one, the total deficit $\sum_{t\in\mathcal{W}}\left(\frac{1}{m}-w_{lt}^{(p)}\right)>0$ must be distributed among the remaining (relevant) features. Hence, for each pair ($p,l)$ there exists at least one relevant feature $v$ such that $w_{lv}^{(p)}>\frac{1}{m}$. ∎
The lemma above guarantees the existence of at least one feature with weight exceeding $\frac{1}{m}$ in some cluster. This motivates the next lemma, which characterises the weight of such features at different values of $p$.
Lemma 3.
Let $w_{lv}^{(p_{0})}\geq\frac{1}{m}+\gamma$ for some $\gamma>0$. Then $w_{lv}^{(p)}>\frac{1}{m}$ for all $p\in(p_{0}-\delta,\,p_{0}+\delta)$, where
$\delta=\gamma\cdot\frac{A(p_{0})^{2}(p_{0}-1)^{2}}{L(p_{0})},$
with
$A(p_{0})=\sum_{u=1}^{m}\left(\frac{D_{lv}^{(p_{0})}}{D_{lu}^{(p_{0})}}\right)^{1/(p_{0}-1)},\quad L(p_{0})=\sum_{u=1}^{m}\left(\frac{D_{lv}^{(p_{0})}}{D_{lu}^{(p_{0})}}\right)^{\frac{1}{p-1})}\left|\log\left(\frac{D_{lv}^{(p_{0})}}{D_{lu}^{(p_{0})}}\right)\right|.$
Proof.
Let $a_{u}=D_{lv}^{(p_{0})}/D_{lu}^{(p_{0})}$. Then
$w_{lv}^{(p)}=\frac{1}{\sum_{u=1}^{m}\left[\frac{D_{lv}}{D_{lu}}\right]^{\frac{1}{p-1}}}=\frac{1}{\sum_{u=1}^{m}a_{u}^{1/(p-1)}}=\frac{1}{A(p)}.$
Hence,
$\frac{d}{dp}w_{lv}^{(p)}=-\frac{A^{\prime}(p)}{A(p)^{2}},\quad\text{where}\quad A^{\prime}(p)=\sum_{u=1}^{m}\frac{a_{u}^{\frac{1}{p-1}}\log a_{u}}{(p-1)^{2}}.$
Define
$L(p)=\sum_{u=1}^{m}a_{u}^{\frac{1}{p-1}}|\log a_{u}|,$
so that
$|A^{\prime}(p)|\leq\frac{L(p)}{(p-1)^{2}},\quad\text{and}\quad|w_{lv}^{\prime}(p)|\leq\frac{L(p)}{A(p)^{2}(p-1)^{2}}.$
Let $\delta$ be the minimum perturbation needed to decrease the weight of $v$ to $1/m$. Then by the Mean Value Theorem,
$\delta=\left(w_{lv}^{(p_{0})}-\frac{1}{m}\right)\cdot\frac{A(p_{0})^{2}(p_{0}-1)^{2}}{L(p_{0})}.$
Since $w_{lv}^{(p_{0})}\geq\frac{1}{m}+\gamma$,
$\delta\geq\gamma\cdot\frac{A(p_{0})^{2}(p_{0}-1)^{2}}{L(p_{0})},$
as claimed. ∎
Building on these lemmas, we now derive sufficient conditions under which a relevant feature is guaranteed to be selected by FS-MWK++. These conditions link the smoothness of feature weights across $p$, the proportion of clusters where the feature appears prominent, and the structure of the $P$ grid. The result formalises the intuitive idea that stability and alignment with cluster structure make a feature likely to stand out.
@@
Let $X$ contain both relevant and noise features, and suppose noise features are drawn independently from a common distribution and are uncorrelated with cluster structure. Let $v$ be a feature such that for a proportion $\alpha$ of the $k$ clusters, there exists a $p_{l}\in P$ satisfying
$w_{lv}^{(p_{l})}\geq\frac{1}{m}+\gamma,$
with $\gamma>0$ and suppose that for all such $p_{l}$,
$\gamma\cdot\frac{A(p_{l})^{2}(p_{l}-1)^{2}}{L(p_{l})}>\frac{1}{2\alpha}.$
Then FS-MWK++ is guaranteed to select $v$, provided that the number of selected features $r$ is at least the number of features satisfying this condition.
@@
Notice there are $|P|\cdot k$ pairs $(p,l)$. To guarantee $v$ is selected, it is sufficient to show that $w_{lv}^{(p_{l})}>\frac{1}{m}+\gamma$ for some $\gamma>0$ in at least $\frac{|P|\cdot k}{2}$ pairs. Let this weight condition hold for $\alpha\cdot k$ clusters for each $p_{l}$ in $\mathcal{P}$. Then,
$\alpha\cdot k\cdot|\mathcal{P}|>\frac{|P|\cdot k}{2}\iff 2\alpha|\mathcal{P}|>|P|\iff|\mathcal{P}|>\frac{|P|}{2\alpha}.$
Recall that $|P|=20$ and that the interval $[p-\delta,p+\delta]$ has length $2\delta$. Given the step within $P$ is of $0.1$, we have
$|\mathcal{P}|\approx\frac{2\delta}{0.1}=20\delta=\delta\cdot|P|>\frac{|P|}{2\alpha}\iff\delta>\frac{1}{2\alpha}.$
Following Lemma 3, the above leads to
$\gamma\cdot\frac{A(p_{l})^{2}(p_{l}-1)^{2}}{L(p_{l})}>\frac{1}{2\alpha}.$
Let us now demonstrate Theorem 1 with an example. Let $m=12$, $\gamma=0.15$, and $\alpha=0.9$. That is, there exists a feature $v$ such that for 90% of the clusters $S_{l}\in S$ there exists a (possibly different) $p_{l}\in P$ for which $w_{lv}^{(p_{l})}\geq\frac{1}{m}+\gamma$. Recall that FS-MWK++ considers a total of $|P|\cdot k$ pairs $(p,l)$. In this example, the feature is stated to have a high weight in only $0.9\cdot k$ pairs (as there is one $p_{l}$ for each cluster). Hence, it is not immediately obvious FS-MWK++ would select $v$ under mild assumptions. Notice that
$w_{lv}^{(p_{l})}=\frac{1}{A(p_{l})}\geq\frac{1}{m}+\gamma\Rightarrow A(p_{l})\leq\frac{1}{\frac{1}{m}+\gamma}=\frac{1}{\frac{1}{12}+0.15}\approx 4.29.$
Let us assume a value of approximately 80%, $A(p_{l})=3.4$, and $p_{l}=2.4$. Recall that $a_{u}=D_{lv}^{(p_{l})}/D_{lu}^{(p_{l})}$, then
$A(p_{l})=\sum_{u=1}^{m}a_{u}^{1/(p_{l}-1)}=\sum_{u=1}^{12}a_{u}^{1/1.4}=3.4,$
leading to, on average, $a_{u}^{1/1.4}\approx 0.283\iff a_{u}\approx 0.171$. Let us use this to approximate $L(p_{l})$,
$L(p_{l})=\sum_{u=1}^{m}a_{u}^{\frac{1}{p-1}}|\log a_{u}|\approx\sum_{u=1}^{12}0.283|\log 0.171|\approx 6.00.$
Then,
$\gamma\cdot\frac{A(p_{l})^{2}(p_{l}-1)^{2}}{L(p_{l})}>\frac{1}{2\alpha}\iff 0.15\frac{3.4^{2}\cdot(2.4-1)^{2}}{6.00}\approx 0.566>\frac{1}{2}.$
Therefore, the condition in Theorem 1 is satisfied, and $v$ is guaranteed to be selected by FS-MWK++.
@@ 4 Experiments setting
We divide our experiments into two main sets to separately evaluate (i) the clustering performance of the proposed MWK++ initialisation method, and (ii) the effectiveness of FS-MWK++ and SFS-MWK++ for unsupervised feature selection. The source code for our experiments can be found at https://github.com/xzhang4-ops1/FSMWK.
@@ 6 Conclusion
Overall, our results highlight the potential of clustering-driven feature weighting as a robust foundation for unsupervised learning tasks. Future work will explore further theoretical guarantees, and extend our methods to more general clustering frameworks beyond MWK++.